# Robust Understanding of Human-Robot Social Interactions through Multimodal Distillation

### v1

@@ 4.6. Key Takeaways
  • The proposed teacher model SocialC3D effectively integrates multimodal inputs (body, face, hand, gaze and raw image) for egocentric social understanding, surpassing SOTAs. Each modality contributes unique and complementary information.
  • The knowledge distillation framework enables the student model to absorb multimodal and clean knowledge from the teacher, improving its ability of capture meaningful information and robustness against spatial-temporal noisy data.
  • The distilled student SocialEgoMobile achieves a trade-off between performance, robustness, and computational cost, which is suitable for deployment on resource-constrained systems.
@@ 5. Conclusion
In this paper, we address the challenges of performance, robustness, and computational efficiency in forecasting users’ social interaction intentions and behaviours from an egocentric perspective on mobile platforms or embedding system. We propose SocialEgoMobile, a lightweight social cue understanding model based on knowledge distillation. In our knowledge distillation framework, the teacher model, SocialC3D, uses multimodal inputs, including body poses, raw images, facial expressions, hand gestures, and gaze patterns, while the student model only relies on deliberate corrupted body pose features. By distilling knowledge from the teacher model’s integrated multimodal representations and downstream task supervision, the student model learns to produce robust predictions under noisy and incomplete input conditions. This work offers an effective and lightweight solution for real-time social understanding, particularly suited for resource-constrained scenarios such as social robots and intelligence social agents. Future work will focus on further enhancing the model’s robustness through real-world deployment and investigating its generalization to complex scenarios, including domain adaptation across viewpoints and environments.
This paper has been submitted to ACM Multimedia 2025.

### v2

@@ 4.6. Key Takeaways
  • SocialC3D surpasses SOTAs by effectively integrating body, face, hand, gaze, and image inputs for egocentric social understanding, demonstrating the value of complementary multimodal cues and the need for socially aware robots capable of comprehensively interpreting human behaviour.
  • The knowledge distillation framework enables the student model to absorb multimodal knowledge from the teacher, improving its ability to capture meaningful information and robustness against spatial-temporal noisy data in a deployment setting.
  • The distilled SocialEgoMobile achieves a great trade-off between performance, robustness, and computational efficiency, which is suitable for deployment on resource-constrained systems, promoting practical and scalable HRI in real-world settings.
@@ 5. Conclusion
In this paper, we address the challenges of performance, robustness, and computational efficiency in forecasting users’ social interaction intentions and behaviours from an egocentric perspective on mobile platforms or embedding systems. We propose SocialEgoMobile, a lightweight social understanding model based on knowledge distillation. In our knowledge distillation framework, the teacher model, SocialC3D uses multimodal inputs, including body poses, raw images, facial expressions, hand gestures, and gaze patterns, while the student model, SocialEgoMobile only relies on deliberately corrupted body pose features. By distilling knowledge from the teacher model’s integrated multimodal social representations and downstream task supervision, the student model learns to produce robust predictions under noisy and incomplete input conditions. This work offers an effective and lightweight solution for real-time social understanding in noisy environments, particularly suited for resource-constrained platforms such as social robots and intelligence social agents. Future work will focus on further enhancing the model’s robustness through real-world deployment and investigating its generalization to complex scenarios, including domain adaptation across viewpoints and environments.